owl shift, Aug 6, 1999

Jim Hirschauer (jhirsch@cuhep.Colorado.EDU)
Fri, 6 Aug 1999 06:00:44 -0600 (MDT)

Shift Summary for owl shift, 6/Aug/99

On Shift: Elliot Cheu, Jim Hirschauer

Hours of beam delivered at all intensities: 7.0
Hours of beam delivered at nominal intensity: 7.0
Hours of E832 data (special runs not included): 6.25
Hours of special runs (laser scans, muon runs...): .25

Accounting of lost time: .5 hrs - starting new run, DDD veto problem

========================================================================
Runs ended this shift: 14125
Run 14125 - 200K - MUON
Run 14127 - in progress - BEAM_832

Tapes brought to FCC this shift: QKC216 - QKC221
QKS216 - QKS221
QKK046 - QKK221
QKB076,QKB077

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Smooth E832 datataking.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
00:20 We reseated DC2 in plane 4, we got heartbeat

14125 00:25 finished muon run, magnet polarity flipped to NEG

14127 00:40 Began run, had ktevana errors which were fixed by
KREND. Errors may be related to recent AZLAT swap

01:35 Laser DYC full. KFREND fixes.

01:55 Onsplit requests QKCC10 to be mounted in ktev4,
tps3d2

05:30 ktevana errors, KFREND

06:40 Elliot noticed that process 10 on plane 1 is no
longer filling the filt histogram. However,there
are 12 processes running. It's possible that the
shared memory writing no longer works on this
process. The next shift should force L3 processes
after the tapes fill up.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

o Force L3 processes after tapes fill up